you are cutting as many 20 1/2 inch pieces from the board to make ladder steps for a tree fort how many steps can you make how much wood is leftover


1
Expert's answer
2021-09-20T16:11:26-0400

"\text{Since we are given that length of pieces from the board to make ladder steps for a tree }\\
\text{fort = $20\frac{1}{2}$ inch = $\frac{41}{2}$inch}\\
\text{Let the length of board = 120inch}\\
\text{N of pieces from the board to make ladder steps for a tree fort is given by}\\
\frac{120}{\frac{41}{2}} = 5\frac{35}{41}\\
\text{Hence the no of steps you can make is 5 steps and $\frac{35}{41}$inch is left over}"